OpenStudy (anonymous):

A square has the side length of (7x-2) in. What is the area of the square?

OpenStudy (anonymous):

Take the side and square it. Area of a square = side^2

OpenStudy (anonymous):

If (7x-2) is one side and we know it's a square then we use the formula length times width to get (7x-2)^2 which to do is 7x*7x = 49x^2 7x*-2 = -14x multiply it by two because we do it twice so -14x*2 = -28x -2*-2 = 4 add them all together and we get 49x^2 - 28x + 4 for the area of the square;) hope i helped! if i did, u should click the best response by my name =D

OpenStudy (anonymous):

You are so bad, you are not supposed to give them answers.

OpenStudy (anonymous):

i did the work too i didn't just give them the answer

OpenStudy (anonymous):

Nice you did the work for the person. /clap

OpenStudy (anonymous):

?

OpenStudy (anonymous):

It's nothing big

myininaya (myininaya):

@kelly226 Please try to engage the asker. Have them do some work. Thanks.

OpenStudy (anonymous):

I agree, but i just wanted to know.

myininaya (myininaya):

We have a code of conduct at www.openstudy.com/code-of-conduct . According to it, we should teach and learn. :)

OpenStudy (anonymous):

well i just made this and i'm already not liking the results of this, so i guess I'm about to delete mine. Have a good day everyone

myininaya (myininaya):

Delete your account?

OpenStudy (anonymous):

yes

myininaya (myininaya):

If you really want your account deleted, I think you can make this request be emailing contact@openstudy.com But can I ask why you don't like about OpenStudy?

myininaya (myininaya):

If you want to pm me, you can.

OpenStudy (anonymous):

Because they already had an argument over someone giving me the answer over something simple and I'm not trying to do that. Why should i already have to block someone and i JUST made this account?

myininaya (myininaya):

We just want to make sure people aren't cheating and are learning. This is the goal of OpenStudy to help educate.

myininaya (myininaya):

You can always post feedback in our feedback group if you don't like something or if you want to comment about something on the site. Would you be interested in doing that? I can give you the link.